A statistics professor plans classes so carefully that the lengths of her classes are uniformly distributed between 46.0 and 56.0 minutes. Find the probability that a given class period runs between 51.25 and 51.75 minutes. 


Find the probability of selecting a class that runs between 51.25 and 51.75 minutes. 
Answer #1

\(X \sim U(46,56)\) is the random variable of the class duration (length) \(P(51.25 \leq X \leq 51.75)=\frac{51.75-51.25}{56-46}=0.05\) is the required probability